OpenStudy (anonymous):

Rationalize the denominator of 9/(7-3√3)-(6-4√3)

OpenStudy (asnaseer):

is it: (9/x) - y OR: 9/(x-y)

OpenStudy (asnaseer):

\[\frac{9}{7-3√3}-(6-4√3)\]or: \[\frac{9}{(7-3√3)-(6-4√3)}\]

OpenStudy (anonymous):

Oh the second one sorry was trying to understand how to do it

OpenStudy (asnaseer):

ok, so this the the expression you want to rationalize, correct? \[\frac{9}{(7-3√3)-(6-4√3)}\]

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

$$ \frac{9}{1+\sqrt{3}} $$

OpenStudy (asnaseer):

ok, first simplify the denominator: \[\frac{9}{(7-3√3)-(6-4√3)}=\frac{9}{7-3√3-6+4√3}=\frac{9}{1+\sqrt{3}}\]

OpenStudy (anonymous):

well now you could rationalize ..

OpenStudy (asnaseer):

you can then remove the radicals from the denominator by noting that:\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

Lol thanks to both of you

OpenStudy (anonymous):

glad to help :)

OpenStudy (asnaseer):

so you can multiply the numerator and denominator by:\[(1-\sqrt{3})\]

OpenStudy (asnaseer):

I assume you are ok from here on?

OpenStudy (anonymous):

Yes thanks again :)

OpenStudy (asnaseer):

yw